OpenStudy (anonymous):

what is the derivative of: (7q)^(1/2) .. and please explain

sam (.sam.):

\[\frac{d}{dq}=\frac{7}{2\sqrt{7}}\]

OpenStudy (agreene):

\[\frac{d}{dq}\sqrt{7q}=\sqrt{7}\frac{d}{dq}\sqrt{q}=\sqrt{7}\frac{1}{2\sqrt{q}}=\frac{\sqrt{7}}{2\sqrt{q}}\] basically, u can take the 7^(1/2) out since it is a constant, and then use the power rule for q^(1/2) and then recombine.

OpenStudy (anonymous):

ok thanks